OpenStudy (anonymous):

Use the technique of completing the square to transform the quadratic equation into the form (x + c)2 = a. 2x2 + 32x + 12 = 0

OpenStudy (anonymous):

\[2x^2+32x+12=0\] to complete the square you need the leading coefficient to be 1, so divide everything by 2 and get \[x^2+16x+6=0\] subtract 6 from both sides to get \[x^2+16x=-6\] and now "complete the square half of 16 is 8 and \(e^2=64\) you write \[(x+8)^2=-6+64=58\]

OpenStudy (anonymous):

i meant "half of \(16\) is \(8\) and \(8^2=64\)"

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

yw